### 07 Jan 2024 v1

# A Generalization of Lévy’s Theorem on Transition Matrices

Abstract.
We generalize a fundamental theorem on transition matrices stating that each component is either strictly positive for all times or identically zero (”Lévy’s Theorem“). Our proof of this fact that does not require the matrices to be Markovian nor to be continuous at time zero. We also provide a formulation of this theorem in the terminology of one-parameter operator semigroups.
Key words and phrases:
transition matrix; operator semigroup; discrete spaces
2020 Mathematics Subject Classification:
Primary 60J35, 47D03; Secondary: 46A40, 47B65
1. Introduction
Let $I$ denote a countable set. A family of functions $p_{ij}\colon(0,\infty)\to[0,\infty)$, $i,j\in I$, is called a transition matrix (over the index set $I$) if for every $i,j\in I$ the following two conditions are fulfilled:
  1. (a)
    $p_{ij}$ is continuous on $(0,\infty)$ and
  2. (b)
    $p_{ij}(t+s)=\sum_{k\in I}p_{ik}(t)p_{kj}(s)$ for all $t,s>0$,
The main result of the present article is the following theorem (see Theorem 2.4):
Theorem 1.1.
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Then for all $i,j\in I$
(1.1) either $\text{either }p_{ij}(t)=0\text{ for all }t>0\text{ or }p_{ij}(t)>0\text{ for all }t>0.$
This statement can directly be transferred to the theory of operator semigroups as follows (see Corollary 3.1 and Example 3.2):
Theorem 1.2.
Let $E$ be one of the sequence spaces $c_{0}$, $c$ or $\ell^{p}$ for some $1\leq p\leq\infty$ (or, more precisely, a discrete Archimedean Riesz in the sense of Corollary 3.1). For $k\in\mathds{N}$ denote by $e_{k}$ the $k$-th canonical unit vectors of $E$. Let $(T_{t})_{t>0}$ be a one-parameter semigroup of positive linear operators on $E$ such that $t\mapsto\langle T_{t}e_{i},e_{j}\rangle$ is continuous on $(0,\infty)$ for all $i,j\in\mathds{N}$ (this is obviously fulfilled whenever the semigroup is strongly continuous).
Then, for every $i,j\in\mathds{N}$, either $\langle T_{t}e_{i},e_{j}\rangle>0$ for all $t>0$ or $\langle T_{t}e_{i},e_{j}\rangle=0$ for all $t>0$.
Under additional assumptions, Theorem 1.1 is well-known in the literature. A first partial proof was given by Lévy in [14], which is why Theorem 1.1 is sometimes referred to as ”Levy’s Theorem“. In [6, Thm 10.1], Chung proved Statement (1.1) in the most general situation so far, where he assumed the transition matrix in addition to be continuous at $t=0$, i.e.
  1. (c)
    $\lim_{t\to 0}p_{ij}(t)=\begin{cases}0&i\neq j\\ 1&i=j\end{cases}$,
Such transition matrices satisfying property (c) are also called standard in the literature. Unfortunately, the proof given in [6, Thm 10.1] is apparently not complete: it is unclear how to conclude $p(2t_{0})\to 0$ as $N\to 0$ from the uniform convergence of (10.2) at the top of page 76, as several objects depend on $N$; however, it is not difficult to understand the author’s intention by taking his previous publications as [7, Thm II.1.5] into account. We also refer to the notes in [7] for further historical references and several different proofs of Theorem 1.1 for transition matrices which are Markovian, i.e.
  1. (d)
    $\sum_{j\in I}p_{ij}(t)=1$ for all $t>0$.
In the theory of operator semigroups a version of Theorem 1.2 for positive analytic $C_{0}$-semigroups on general (not necessarily discrete) Banach lattices can be found in [2, Thm C-III 3.2(b)]. For self-adjoint $C_{0}$-semigroups on $L^{2}$-spaces this result is due to Simon [16, Thm 1]. For $C_{0}$-semigroups on $L^{p}$-spaces it was shown by Kishimoto and Robinson [11, Lem 8] and it was generalized to the Banach lattice setting by Majewski and Robinson [15, Prop 2]. Jochen Glück pointed out to the author that, similar to the main result of the present article, the assumption of strong continuity at time zero can also be dropped in the analytic case. This result is provided as Theorem A.1 in the appendix. In the 21th century, results similar to Theorem 1.2 have also been proven in the context of infinite graphs, cf. [9, Thm 7.3] and [10, Thm 1.26], which is a special case of both, the discrete and the analytic setting.
The present article has three goals: First, to give a complete proof of Chung’s result. Second, to generalize it by dropping both assumptions (c) and (d). And third, to transfer the result to the theory of operator semigroups, where it is apparently little known.
2. Transition Matrices
The proof of the following theorem is a combination and adaption of [7, Thm II.1.5] and [6, Thm 10.1].
Theorem 2.1.
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$. Then $p_{ij}(t)=0$ for all $0<t\leq t_{0}$ implies $p_{ij}(t)=0$ for all $t>0$.
Proof.
Aiming for a contradiction, first assume that $c\coloneqq p_{ij}(2t_{0})>0$. Fix $70<\varepsilon<c/$ 7 and define $\delta\coloneqq 1/16$. By Dini’s theorem, the series
$c=p_{ij}(2t_{0})=\sum_{k\in I}p_{ik}(t)p_{kj}(2t_{0}-t)$
converges uniformly on $[\delta 2t_{0},(1-\delta)2t_{0}]$. Hence, there exists a finite set $F\subseteq I$ such that
$\sum_{k\in I\setminus F}p_{ik}(t)p_{kj}(2t_{0}-t)<\varepsilon\text{ for all }t\in[\delta 2t_{0},(1-\delta)2t_{0}].$
Let $N\coloneqq\#F$ denote the number of elements in $F$. By enlarging the set $F$ if necessary we may assume that $N$ is an even number. We define the stepwidth $s\coloneqq\frac{t_{0}}{4N}$ and the abbreviations $N_{0}\coloneqq 8\delta N$ and $N_{1}\coloneqq 8(1-\delta)N$. Note that $N_{0},N_{1}\in\mathds{N}$, $N_{0}s=\delta 2t_{0}$, $N_{1}s=(1-\delta)2t_{0}$ and $(N_{0}+N_{1})s=2t_{0}$.
Next, for all $m\in\mathds{N}$ we define the sets $C_{m}\coloneqq\{k\in I:p_{ik}(ls)=0\text{ for all }1\leq l\leq m\}$. Since $C_{m+1}\subseteq C_{m}$ for all $m\in\mathds{N}$, the sets $D_{m}\coloneqq C_{m}\setminus C_{m+1}$ are pairwise disjoint. While $j\in C_{m}$ for all $1\leq m\leq 4N$, other sets $C_{m}$ are possibly empty. For $m,n\in\mathds{N}$, $n<8N$, we define
$u(m,n)\coloneqq\sum_{k\in C_{m}}p_{ik}(ns)p_{kj}(8Ns-ns)\text{ and }v(m,n)\coloneqq\sum_{k\in D_{m}}p_{ik}(ns)p_{kj}(8Ns-ns).$
By definition of $C_{m}$, $u(m+1,n)\leq u(m,n)$ for all $m,n$ and $u(m,n)=0$ whenever $n\leq m$. Moreover, $v(m,n)<\varepsilon$ whenever $D_{m}$ is disjoint from $F$ and $n\in\{N_{0},\dots,N_{1}\}$. Given $M,m\in\mathds{N}$, $k\in C_{M}$ and $l\in I$ such that $p_{lk}(ms)>0$, then $l\in C_{M-m}$. In fact, otherwise it would follow from
$p_{ik}(M^{\prime}s)\geq p_{il}\big{(}(M^{\prime}-m)s\big{)}p_{lk}(ms)>0\text{ for some }m+1\leq M^{\prime}\leq M$
that $k\not\in C_{M}$. From this, we now conclude the following two assertions: First,
$u(m,n+1)-v(m,n+1)$ $=\sum_{k\in C_{m+1}}\biggl{(}\sum_{l\in I}p_{il}(ns)p_{lk}(s)\biggr{)}p_{kj}(8Ns-ns-s)$
$=\sum_{k\in C_{m+1}}\biggl{(}\sum_{l\in C_{m}}p_{il}(ns)p_{lk}(s)\biggr{)}p_{kj}(8Ns-ns-s)$
$\leq\sum_{l\in C_{m}}p_{il}(ns)\sum_{k\in I}p_{lk}(s)p_{kj}(8Ns-ns-s)=u(m,n)$
for all $m,n\in\mathds{N}$, $n+1<8N$; and second, as $j\in C_{4N}$,
$c=p_{ij}(2t_{0})=\sum_{l\in I}p_{il}(N_{1}s)p_{lj}(N_{0}s)=\sum_{l\in C_{4N-N_{0}}}p_{il}(N_{1}s)p_{lj}(N_{0}s)=u(4N-N_{0},N_{1}).$
Combining both observations, we obtain for all $N_{0}\leq m\leq 4N-N_{0}$
(2.1) $c=u(4N-N_{0},N_{1})\leq u(m,N_{1})=u(m,N_{1})-u(m,N_{0})\leq\sum_{n=N_{0}+1}^{N_{1}}v(m,n).$
Recall that at most $N$ of the disjoint sets $D_{m}$ can have a non-empty intersection with $F$. As the set $\{N_{0},\dots,4N-N_{0}\}$ of all choice for $m$ has $4N-2N_{0}+1\geq(4-16\delta)N=3N$ elements, we find $M\subseteq\{N_{0},\dots,4N-N_{0}\}$ such that #$\#M=N$ and $D_{m}\subseteq I\setminus F$ for all $m\in M$. Summing over $m\in M$, we obtain from (2.1)
$Nc$ $\leq\sum_{m\in M}\sum_{n=N_{0}+1}^{N_{1}}v(m,n)=\sum_{n=N_{0}+1}^{N_{1}}\sum_{m\in M}v(m,n)$
$\leq\sum_{n=N_{0}+1}^{N_{1}}\sum_{k\in I\setminus F}p_{ik}(ns)p_{kj}(8Ns-ns)$
$<\sum_{n=N_{0}+1}^{N_{1}}\varepsilon<(N_{1}-N_{0})\varepsilon=(8N-16\delta N)\varepsilon=7N\varepsilon<Nc,$
a contradiction. This proves that $p_{ij}(2t_{0})=0$.
Applying this argument to every smaller value of $t_{0}$ shows that $p_{ij}(2t)=0$ for every $0<t\leq t_{0}$. Finally, we obtain by induction that $p_{ij}(t)=0$ for all $t>0$. ∎
Corollary 2.2.
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t\geq t_{0}$
Proof.
If $p_{ij}(t)=0$ for all $0<t<t_{0}$, then the assertion follows from Theorem 2.1. Otherwise, there exists $0<t_{1}<t_{0}$ such that $p_{ij}(t_{1})>0$. Let $t^{\prime}_{0}\coloneqq\inf\{t>t_{1}:p_{ij}(t)=0\}$. Since $t^{\prime}_{0}\leq t_{0}$ and $p_{ij}(t^{\prime}_{0})=0$ due to the continuity of $p_{ij}$, we may assume without loss of generality that $t_{0}$ equals $t^{\prime}_{0}$. Now we consider
$I_{1}\coloneqq\{k\in I:\text{ there exists }0<\delta<t_{0}\text{ such that }p_{ik}(t)>0\text{ for all }t_{0}-\delta<t<t_{0}\}$
and define $I_{0}\coloneqq I\setminus I_{1}$. Note that for every $k\in I_{0}$ it follows from the continuity of $p_{ik}$ that $p_{ik}(t_{0})=0$. Let $k\in I_{1}$ and let $0<\delta<t_{0}$ such that $p_{ik}(t)>0$ for all $t_{0}-\delta<t<t_{0}$. For every $0<s<\delta$ it follows from
$0=p_{ij}(t_{0})\geq p_{ik}(t_{0}-s)p_{kj}(s)$
that $p_{kj}(s)=0$. Consequently, by Theorem 2.1, $p_{kj}(t)=0$ for all $t>0$. Combining $p_{ik}(t_{0})=0$ for $k\in I_{0}$ and $p_{kj}(t)=0$ for all $t>0$ and $k\in I_{1}$, we obtain that
$p_{ij}(t_{0}+t)=\sum_{k\in I_{0}}p_{ik}(t_{0})p_{kj}(t)+\sum_{k\in I_{1}}p_{ik}(t_{0})p_{kj}(t)=0$
for all $t>0$. This completes the proof. ∎
In order to obtain Theorem 1.1, it remains to show that $p_{ij}(t_{1})>0$ implies that $p_{ij}(t)>0$ for all $t\geq t_{1}$; a fact that is easy to see for Markovian (or standard) transition matrices, cf. [7, Thm II.1.4.], but surprisingly sophisticated to achieve in the general setting of this article. This is the aim of what follows.
Lemma 2.3.
Let $p_{ij}$ be a transition matrix over a countable index set $I$ and let $j\in I$ and assume that $p_{jj}(t)=0$ for all $t>0$.
  1. (a)
    Let $t>0$, $i,k\in I$ and $(\alpha_{n})\subseteq(0,\infty)$. Then for every $N\in\mathds{N}$ one has
    (2.2) $p_{ik}\biggl{(}t+\sum_{n=1}^{N}\alpha_{n}\biggr{)}=\sum_{n=1}^{N}p_{ij}\biggl{(}t+\sum_{m=1}^{n-1}\alpha_{m}\biggr{)}p_{jk}\biggl{(}\sum_{m=n}^{N}\alpha_{m}\biggr{)}$
    $\quad+\sum_{k_{1},\dots,k_{N}\neq j}p_{ik_{1}}(t)p_{k_{1}k_{2}}(\alpha_{1})\dots p_{k_{N-1}k_{N}}(\alpha_{N-1})p_{k_{N}k}(\alpha_{N}).$
  2. (b)
    Let $i\in I$ and $t_{1}>0$ such that $p_{ij}(t_{1})>0$. Then $p_{jk}(t)=0$ for all $k\in I$ and $t>0$.
Proof.
(a) We prove the statement by induction over $N$. For $N=1$ one has
$p_{ik}(t+\alpha_{1})=\sum_{k_{1}\in I}p_{ik_{1}}(t)p_{k_{1}k}(\alpha_{1})=p_{ij}(t)p_{jk}(\alpha_{1})+\sum_{k_{1}\neq j}p_{ik_{1}}(t)p_{k_{1}k}(\alpha_{1}),$
which corresponds to (2.2). Now assume that (2.2) holds for a given $N\in\mathds{N}$. Using that $p_{jj}$ is identically zero, this implies for $k=j$ that
(2.3) $\sum_{k_{1},\dots,k_{N}\neq j}p_{ik_{1}}(t)p_{k_{1}k_{2}}(\alpha_{1})\dots p_{k_{N-1}k_{N}}(\alpha_{N-1})p_{k_{N}j}(\alpha_{N})$ $=p_{ij}\biggl{(}t+\sum_{n=1}^{N}\alpha_{n}\biggr{)}.$
Now it follows that
$p_{ik}\biggl{(}t+\sum_{n=1}^{N+1}\alpha_{n}\biggr{)}=\sum_{k_{N+1}\in I}p_{ik_{N+1}}\biggl{(}t+\sum_{n=1}^{N}\alpha_{n}\biggr{)}p_{k_{N+1}k}(\alpha_{N+1})$
$=\sum_{k_{N+1}\in I}\biggl{(}\sum_{n=1}^{N}p_{ij}\biggl{(}t+\sum_{m=1}^{n-1}\alpha_{m}\biggr{)}p_{jk_{N+1}}\biggl{(}\sum_{m=n}^{N}\alpha_{m}\biggr{)}$
$\quad+\sum_{k_{1},\dots,k_{N}\neq j}p_{ik_{1}}(t)p_{k_{1}k_{2}}(\alpha_{1})\dots p_{k_{N-1}k_{N}}(\alpha_{N-1})p_{k_{N}k_{N+1}}(\alpha_{N})\biggr{)}p_{k_{N+1}k}(\alpha_{N+1})$
$=\sum_{n=1}^{N}p_{ij}\biggl{(}t+\sum_{m=1}^{n-1}\alpha_{m}\biggr{)}p_{jk}\biggl{(}\sum_{m=n}^{N+1}\alpha_{m}\biggr{)}+p_{ij}\biggl{(}t+\sum_{m=1}^{N}\alpha_{m}\biggr{)}p_{jk}(\alpha_{N+1})$
$\quad+\sum_{k_{1},\dots,k_{N+1}\neq j}p_{ik_{1}}(t)p_{k_{1}k_{2}}(\alpha_{1})\dots p_{k_{N}k_{N+1}}(\alpha_{N})p_{k_{N+1}k}(\alpha_{N+1}),$
where the second equation follows from induction hypothesis (2.2) for $k=k_{N+1}$ and the third from (2.3) and from $p_{jj}=0$. This shows that (2.2) holds for $N+1$.
(b) Aiming for a contradiction, assume that $p_{jk}(t_{2})>0$ for some $t_{2}>0$ and $k\in I$. By continuity, there exists $\varepsilon>0$ and $\delta>0$ such that $p_{ij}(t_{1}+t)\geq\varepsilon$ and $p_{jk}(t_{2}+t)\geq\varepsilon$ for all $t\in[0,\delta]$. Choose a sequence $(\alpha_{n})\subseteq(0,\infty)$ such that $\sum_{n\in\mathds{N}}\alpha_{n}<\delta$. For every $N\in\mathds{N}$, it now follows from Part (a) that
$\sup_{t\in[0,\delta]}p_{ik}(t_{1}+t+t_{2})$ $\geq p_{ik}\biggl{(}t_{1}+\sum_{n=1}^{N}\alpha_{n}+t_{2}\biggr{)}$
$\geq\sum_{n=1}^{N+1}p_{ij}\biggl{(}t_{1}+\sum_{m=1}^{n-1}\alpha_{m}\biggr{)}p_{jk}\biggl{(}\sum_{m=n}^{N}\alpha_{m}+t_{2}\biggr{)}$
$\geq\sum_{n=1}^{N+1}\varepsilon^{2}=(N+1)\varepsilon^{2}.$
This leads to a contradiction as $N\to\infty$. Thus, $p_{jk}(t)=0$ for all $t>0$. ∎
The assumption $p_{ij}(t_{1})>0$ for some $t_{1}>0$ in Part (b) if Lemma 2.3 may look a bit artifical at first glance. However, the example
$p_{ij}=\begin{cases}1&j=k\\ 0&j\neq k\end{cases}$
of a constant transition matrix shows that this assumption is indeed necessary. The next theorem contains the article’s main result.
Theorem 2.4.
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t>0$.
Proof.
We prove that for every $k\in I$ at least one of the functions $p_{ik}$ and $p_{kj}$ is identically zero. To this end, let $k\in I$ and suppose that $p_{ik}(v)>0$ for some $v>0$. By Corollary 2.2, $p_{ik}(t)>0$ for all $0<t<v$. Hence, we may assume without loss of generality that $v<t_{0}$. It then follows from
$0=p_{ij}(t_{0})\geq p_{ik}(v)p_{kj}(t_{0}-v)$
that $p_{kj}(t_{0}-v)=0$. Aiming for a contradiction, we suppose that $p_{kj}$ is not identically zero. In this case, by Corollary 2.2,
(2.4) $u\coloneqq\inf\{t>0:p_{kj}(t)=0\}>0$
and for each $0<s<u$ it follows from
$0=p_{kj}(u)\geq p_{kk}(s)p_{kj}(u-s)$
and $p_{kj}(u-s)>0$ that $p_{kk}(s)=0$. Now Theorem 2.1 implies that $p_{kk}(t)=0$ for all $t>0$. Theorefore, $p_{kl}(t)=0$ for all $t>0$ und $l\in I$ by Part (b) of Lemma 2.3. In paticular, $p_{kj}$ is identically zero in contradiction to our assumption.
Hence, for every $t>0$ one has
$p_{ij}(t)=\sum_{k\in I}p_{ik}(t/2)p_{kj}(t/2)=0.$
This completes the proof. ∎
3. Operator Semigroups on Discrete Spaces
In what follows, we give a reformulation of Theorem 2.4 for one-parameter semigroups on discrete Riesz spaces. Recall that a family of linear operators $(T_{t})_{t\in(0,\infty)}$ on a vector space $X$ is called a (one-parameter operator) semigroup if $T_{t+s}=T_{t}T_{s}$ for all $t,s>0$.
One-parameter operator semigroups appear naturally to describe solutions of linear autonomous evolution equations. In many applications the underlying Banach space is a function space and thus exhibits some kind of order structure. If, in such a situation, a positive initial value of the evolution equation leads to a positive solution, one speaks of a positive operator semigroup. We refer to the recent monograph [5] for a modern introduction to the theory of positive semigroups.
We recall some basic notations from the theory of Riesz spaces according to [13]. A real ordered vector space is called a Riesz space (or vector lattice) if for every pair $u,v\in E$ the supremum $u\vee v\coloneqq\sup\{u,v\}$ exists in $E$. In this case, one defines $\lvert u\rvert\coloneqq u\vee(-u)$ and $u\wedge v\coloneqq-\big{(}(-u)\vee(-v)\big{)}$ for all $u,v\in E$ and $E_{+}\coloneqq\{u\in E:u\geq 0\}$. Two vectors $u,v\in E$ are called disjoint if $\lvert u\rvert\wedge\lvert v\rvert=0$. A linear operator $T\colon E\to F$ between Riesz spaces $E$ and $F$ is called positive if $TE_{+}\subseteq F_{+}$; an operator semigroup is called positive if it consists of positive operators. A Riesz space $E$ is called Archimedean if for every $u\in E_{+}$ such that the set $\{nu:n\in\mathds{N}\}$ is order bounded in $E$ it follows that $u=0$.
Let $E$ be an Archimedean Riesz space. A vector subspace $J\subseteq E$ such that $\lvert u\rvert\leq\lvert v\rvert$ for $v\in J$ implies that $u\in J$ is called an ideal. The smallest ideal containing a given $u\in E_{+}$ is called the principal ideal generated by $u$ and is given by $E_{u}=\{v\in E:\exists c\geq 0\text{, }\lvert v\rvert\leq cu\}$. If an ideal $J\subset E$ is closed under general suprema, i.e., $\sup A\in J$ for each set $A\subseteq J$ whose supremum exists in $E$, then $J$ is called a band. For each subset $A\subseteq E$ we denote by $A^{\bot}\coloneqq\{v\in E:\lvert v\rvert\wedge\lvert u\rvert=0\text{ for all }u\in A\}$ the disjoint complement of $A$; we recall that $A^{\bot}$ is a band and that $A^{\bot\bot}\coloneqq(A^{\bot})^{\bot}$ is the smallest band containing $A$. A vector $u\in E_{+}$ is called a weak (order) unit (of $E_{+}$) if the principal band $\{u\}^{\bot\bot}$ is equal to the whole space $E$.
An element $d\in E$, $d\neq 0$, is called discrete if the principal ideal $E_{d}$ is one-dimensional, i.e. $E_{d}=\{\lambda d:\lambda\in\mathds{R}\}$. Since $E$ is Archimedean, the principal band generated by a discrete element $d$ is given by $\{d\}^{\bot\bot}=\{\lambda d:\lambda\in\mathds{R}\}$ and is a projection band, i.e. $\{d\}^{\bot\bot}$ is the range of a positive projection $P_{d}\colon E\to\{d\}^{\bot\bot}$ such that $I-P_{d}$ is a positive projection onto $\{d\}^{\bot}$ [13, Thm 26.4]. Since the band projection is uniquely determined, for every discrete element $d\in E$ there exists a uniquely determined linear functional $\varphi_{d}\colon E\to\mathds{R}$ such that $\langle\varphi_{d},u\rangle d=P_{d}u$ for all $u\in E$. Clearly, $\varphi_{d}$ is positive if and only if $d\in E_{+}$. We call the Archimedean Riesz space $E$ discrete if there exists a set $D\subseteq E$ of discrete elements such that $D^{\bot\bot}=E$.
Using this terminology, Theorem 2.4 can be reformulated as follows.
Corollary 3.1.
Let $E$ be a discrete Archimedean Riesz space and $D\subseteq E_{+}$ be a countable set of pairwise disjoint discrete elements such that $D^{\bot\bot}=E$. Let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on $E$ such that $t\mapsto\langle\varphi_{d},T_{t}d^{\prime}\rangle$ is continuous on $(0,\infty)$ for all $d,d^{\prime}\in D$.
Then for all $d\in D$, either $T_{t}d$ is weak order unit for all $t>0$ or $T_{t}d=0$ for all $t>0$.
Example 3.2.
Let $E$ denote one of the following sequence spaces: $c$ (space of convergent sequences), $c_{0}$ (the space of nullsequences), $l^{p}$ (space of $p$-summable sequences for $1\leq p<\infty$) or $\ell^{\infty}$ (space of bounded sequences). Let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup that is strongly continuous, i.e. $t\mapsto{T_{t}x}$ is continuous on $(0,\infty)$ for all $x\in E$. Then for all $x\in E_{+}$, either $T_{t}x$ is a weak order unit for all $t>0$ or $T_{t}x=0$ for all $t>0$.
Remark 3.3.
While Corollary 3.1 follows immediately from Theorem 2.4 applied to the transition matrices $p_{ij}(t)=\langle T_{t}d_{i},\varphi_{d_{j}}\rangle$ (for discrete elements $d_{i},d_{j}\in D$), it is also possible to conclude Theorem 2.4 from the Corollary as follows: Let $p_{ij}$ be a transition matrix over the index set $\mathds{N}$. On $\ell^{0}$, the discrete Archimedean Riesz space of all real sequences, we consider the set
$E\coloneqq\{(x_{n})\in\ell^{0}:\sum\nolimits_{n}\lvert x_{n}\rvert p_{nm}(t)<\infty\text{ for all }t>0\text{ and }m\in\mathds{N}\},$
which is an ideal in $\ell^{0}$ and hence itself a discrete Archimedean Riesz space that contains the space $c_{00}$ of all finite sequences. On this space, we define the mappings $T_{t}(x_{n})\coloneqq\big{(}\sum_{n}x_{n}p_{nm}(t)\big{)}_{m}$. By definition of $E$, $T_{t}\colon E\to E$ is an one-parameter semigroup that satisfies all conditions of Corollary 3.1. Hence, for every $n,m\in\mathds{N}$, $p_{nm}(t)=\langle T_{t}e_{n},e_{m}\rangle$ is either strictly positive for all $t>0$ or identically zero.
Appendix A Analytic Semigroups
In what follows, a one-parameter operator semigroup $(T_{t})_{t\in(0,\infty)}$ on a Banach space $X$ is called analytic (or holomorphic) if there exists $\theta\in(0,\frac{\pi}{2}]$ such the operator valued mapping $t\mapsto T_{t}$ has an analytic extension to $\Sigma_{\theta}\coloneqq\{z\in\mathds{C}\setminus\{0\}:\lvert\arg z\rvert<\theta\}$. In the literature the definition of analyticity typically requires in addition a boundedness condition in the neighborhood of the origin, cf. [1, Def 3.7.1] or [12, Ch 2], which is not needed in this section. As the most important example, we mention that on a Hilbert space every semigroup generated by a self-adjoint operator that is bounded above is analytic, cf. [8, Cor II.4.7]. For the sake of completeness, we recall that a semigroup $(T_{t})_{t\in(0,\infty)}$ is called a $C_{0}$-semigroup if $\lim_{t\downarrow 0}T_{t}x=x$ for all $x\in X$. An interesting example of an analytic semigroup that is not $C_{0}$ is for instance given by heat euqation with non-local boundary conditions on $L^{\infty}(\Omega)$, cf. [3, Thm 4.8].
Theorem A.1.
Let $E$ be a Banach lattice and let $(T_{t})_{t\in(0,\infty)}$ be an analytic positive semigroup on $E$. Let $f\in E_{+}$ and $\varphi\in E^{\prime}_{+}$. Then either $\langle\varphi,T_{t}f\rangle=0$ for all $t\in(0,\infty)$ or $\langle\varphi,T_{t}f\rangle>0$ for all $t\in(0,\infty)$.
A result in the same vein was proven in [4, Prop 3.12] for eventually positive rather than positive semigroups; while the semigroups considered there are $C_{0}$, the fact that eventual positivity suffices indicates that the behaviour at time 00 is not important. The proof of Thm A.1 is an adaptation of the classical proof for $C_{0}$-semigroups and was communicated to the author by Jochen Glück.
The argument mainly follows the somewhat simplified presentation of the $C_{0}$-semigroup proof from [2, Thm C-III 3.2(b)] – yet, the argument becomes still a bit more transparent if one splits off the following observation, that does not need analyticity of the semigroups, into a separate lemma.
Lemma A.2.
Let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on a Banach lattice $E$. Let $f\in E_{+}$ and assume that $T_{t}f\to f$ for $t\downarrow 0$. Then there exists a sequence $t_{n}\downarrow 0$ and an increasing sequence $(f_{n})$ in $E$ that converges to $f$ and that satisfies $0\leq f_{n}\leq T_{t_{n}}f$ for all $n$.
Proof.
Let $(t_{n})$ converge to 00 sufficiently fast that $\sum_{n}\lVert f-T_{t_{n}}f\rVert<\infty$ and define
$\tilde{f}_{n}\coloneqq f-\sum_{k=n}^{\infty}\big{(}f-T_{t_{k}}f\big{)}^{+}$
for each $n$. Then $(\tilde{f}_{n})$ is an increasing sequence in $E$ that converges to $f$ and for each $n$ one has $\tilde{f}_{n}\leq f-(f-T_{t_{n}}f)^{+}\leq T_{t_{n}}f$. By defining $f_{n}\coloneqq\tilde{f}_{n}^{+}$ for each $n$ one obtains a sequence $(f_{n})$ in $E_{+}$ with the desired properties. ∎
Proof of Theorem A.1.
Let $\langle\varphi,T_{s}f\rangle=0$ for some time $s>0$. Step 1: Assume, in addition, that $T_{t}f\downarrow f$ as $t\downarrow 0$. Then we can choose sequences $(t_{n})$ in $(0,s)$ and $(f_{n})$ in $E_{+}$ as in Lemma A.2. For each fixed index $m$ and all $n\geq m$ we get
$0\leq T_{s-t_{n}}f_{m}\leq T_{s-t_{n}}f_{n}\leq T_{s}f$
so $\langle\varphi,T_{s-t_{n}}f_{m}\rangle=0$ for all $n\geq m$. Since the sequence $(s-t_{n})$ clusters at $s$, the identity theorem for analytic functions implies that $\langle\varphi,T_{t}f_{m}\rangle=0$ for all $t\in(0,\infty)$. By letting $m\to\infty$ we conclude that $\langle\varphi,T_{t}f\rangle=0$ for all $t\in(0,\infty)$. Step 2: Now let $f$ be general. Due to the analyticity of the semigroup the orbit of $f$ is continuous on $(0,\infty)$, so Step 1 can be applied to the vector $\tilde{f}:=T_{s/2}f$ since $\langle\varphi,T_{s/2}\tilde{f}\rangle=0$. So $\langle\varphi,T_{t}f\rangle=0$ for all $t>s/2$ and another application of the identity theorem shows that the same equality remains true for all $t>0$. ∎
Step 1 in the proof above is simply the argument for the $C_{0}$-semigroup case. The main point is to observe that this argument does not really need the entire semigroup to be $C_{0}$, but just the orbit of $f$.